OpenStudy (anonymous):

HELP ME PLZ A new computer costs $899.00 at Computer World. You can pay that amount in cash, or you can choose a credit plan and pay a $180.00 down payment and then monthly installments of $40.00 for 2 years. How much more will you pay for the computer if you choose the credit plan?

OpenStudy (tkhunny):

Have you considered adding up the payments?

OpenStudy (anonymous):

1,119

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

2,238

OpenStudy (tkhunny):

Let's see... 180 + 24*40 = 180 + 960 = 1140

OpenStudy (anonymous):

ok

OpenStudy (tkhunny):

Now answer the question.

OpenStudy (anonymous):

so 180+24*40=1,140

OpenStudy (anonymous):

180+960=1,140

OpenStudy (anonymous):

then do you divided and get 1

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

You must read and understand the question. How Much More? 1140 - 899 = ??